3x-4y=-19
2x+ 4y=14

Answer by mananth(16949)

3 x -4 y = -19 .............1
2 x + 4 y = 14 .............2

multiply (1)by 1
Multiply (2) by 1
3 x -4 y = -19
2 x + 4 y = 14
Add the two equations
5 x = -5
/ 5
x = -1
plug value of x in (1)
3 x -4 y = -19
-3 -4 y = -19
-4 y = -19 + 3
-4 y = -16
y = 4